### v1

@@ 1.5 Angular Momentum of Bondi-Hoyle Infall in the Simulation
The simulation can be used also to measure $j_{\rm BH}$ relative to the position and velocity of PMS stars, and to compare the result with the prediction of Equation 4. Although new stars are continuously formed in the simulation, towards the end of the run a significant fraction of them have ages in the approximate range 0.5-2.0 Myr, old enough to be representative of Class II PMS stars. However, because the time it takes to assemble a star may vary from star to star, and can be relatively long ($\sim 1$ Myr) for massive stars [61, 62], we select PMS stars based on the local gas density, rather than the stellar age, which better reflects the observational SED classification as well. For that purpose, we use seven snapshots at regular time intervals covering the last 0.75 Myr of the simulation, yielding a total of 1,629 stellar positions.
@@
Averaging over all seven snapshots, we find that the gas density sampled in spheres of radius $\sim 400$ AU (of the order of the size of the largest observed disks) centered around the stars, $P_{\rm st}(n)$, has a clear bimodal distribution, shown by the shaded blue histogram in the left panel of Figure 6. The red unshaded histogram shows the overall gas density distribution sampled uniformly in the whole volume at the same resolution, $P_{\rm V}(n)$. The lower-density peak of $P_{\rm st}(n)$ follows approximately the overall distribution, though shifted to slightly larger density and with a shape a bit skewed to the right, while the higher-density peak has a maximum at a density larger by four orders of magnitude. The stars at such high local density are still embedded in their native dense gas and are generally increasing their mass at a high rate. The black circles in the left panel of Figure 6 show the gas infall rate on the sink particles in the simulation (with values shown on the right $y$ axis). It strongly correlates with the local gas density, and the infall rates of stars below $\sim 10^{5}$ cm-3 are of the same order of magnitude as the observed accretion rates of young PMS stars.
Figure 6: Left: Probability distribution of density sampled in spheres of radius $\sim 403$ au centered around the stars, $P_{\rm st}(n_{\rm H})$ (blue shaded histogram), and sampled uniformly in the whole volume at the same resolution, $P_{\rm V}(n_{\rm H})$ (red unshaded histogram), using all seven time snapshots of the simulation. The black points give the infall rate on the stars from the simulation, averaged over a period of 5,000 yr, with values shown in the right $y$ axis and a linear least-squares fit shown by the long dashed line. The vertical dashed line corresponds to the critical density, $n_{\rm H,cr}=5\times 10^{4}$ cm-3, used to select the stars representative of Class II objects in the BH phase. Right: The predicted BH infall rate versus the infall rate of the stars in the simulation (as in the left plot). The red symbols are the Class II stars based on the critical density criterion, and the red line is the least-squares fit to the red symbols. The black line is the one-to-one relation.
@@
We have verified through visualizations that most of these stars at lower densities are accompanied by gas structures with morphology and kinematics consistent with BH trails, and their predicted BH infall rate is also of the order of the measured infall rate in the simulation, as shown in the right panel of Figure 6. We have also verified that, on average, the local density is inversely correlated with the stellar age. The stars at low density have clearly decoupled from their native dense cores and tend to sample the random density and velocity fields of the parent cloud at larger scales. Based on these results, we choose a fixed critical density value, $n_{\rm H,cr}$, to select the sink particles representative of Class II PMS stars in the BH phase, $n_{\rm H,cr}=5\times 10^{4}$ cm-3, shown by the vertical dashed line in Figure 6. This selection yields 961 PMS stars, out of the total 1,629 stars found in the seven snapshots.
@@
We measure the mean mass, gas velocity and angular momentum within spheres of different radii centered on the position of each star. The density and velocity fields of the whole 4 pc box are first extracted into a uniform grid of 1,0243 cells, so the cell size is $dx=0.004$ pc or 780 AU. The spheres have radii of 2, 4, 8, 16, 32, 64, 128, and 256 $dx$, with the two smallest ones and the largest one expected to be outside of the inertial range of the simulation. Gas mass, and velocity and angular momentum components are averaged within each sphere with a Gaussian cutoff, as in Equation 11.
@@
We estimate the angular momentum of the infalling gas by computing $j$ within a sphere of radius equal to $R_{\rm BH}$, so we need to compute $R_{\rm BH}$ as defined in Equation 3. The mass and velocity of each star are known, as well as the isothermal sound speed in the simulation, $c_{\rm s}=0.18\times 10^{5}$ cm s-1. To compute $v_{\rm rel}$, the difference between the star velocity and the gas velocity, we use the gas mean velocity measured within a sphere centered on the star. Because the flow is turbulent, the mean gas velocity may vary when measured at different scales, so we should compute it at a scale larger than $R_{\rm BH}$, but not too large. We settle on a radius of 8 $dx$=6,250 AU, which is significantly larger than the largest values of $R_{\rm BH}$ and the typical PD sizes. For the sink particles, we find values of $R_{\rm BH}$ in the approximate range 1-3,000 AU. We cannot measure $j$ directly from the simulation at such scales, because they are well within the numerical dissipation range. Instead, we measure it within spheres of radius of 8 $dx$=6,250 AU, as for the estimate of $v_{\rm rel}$, because this is the smallest size we can consider without being significantly affected by numerical dissipation. We then extrapolate the value of $j$ at $R=R_{\rm BH}$, using the linear $j$ scaling established earlier both analytically and numerically (see Figure 5).
@@
The results are shown in Figure 2 as a function of the stellar mass for all the selected PMS stars in the simulation (blue dots). The least-squares fit has a slope of 0.86 (solid blue line), a bit shallower than the one predicted in Equation 8. Figure 2 shows a significant scatter at any given value of $M_{\rm star}$, as expected from Equation 4 due to the dependence on $\sigma_{v,\rm rel}$: for an individual star at a given time, $j_{\rm BH}$ depends on the local value of $v_{\rm rel}$, which is a random variable. The scatter and the average value of $j$ in PDs are expected to be somewhat smaller than those in individual BH spheres computed here, because of partial cancellation in the vector sums of $\bm{J}$ from different BH spheres. Figure 2 shows that observed disks have indeed a slightly reduced mean angular momentum and a smaller scatter than our individual BH spheres, as expected.
@@ 1.6 Time Dependence of PMS Disk Formation by Bondi-Hoyle Infall
Whether or not BH infall is so dominant to completely restructure PDs during the Class II phase, that is on a timescale of $\sim 2$ Myr, depends on the stellar mass and on the time evolution of the relative velocity between the stars and the turbulent gas, $v_{\rm rel}$. The BH infall rate, defined as a mass flux through a surface of area $\pi R_{\rm BH}^{2}$ with gas velocity $v=(c_{\rm s}^{2}+v_{\rm rel}^{2})^{1/2}$, gas number density $n_{\rm H}$, and $R_{\rm BH}$ given by Equation 3 is:
@@
When the stars are fully decoupled from the parent gas, their velocities are not correlated to the gas velocity anymore, and we can assume that the rms velocities satisfy the relation $\sigma^{2}_{v,\rm rel}=\sigma^{2}_{v,0}+\sigma^{2}_{v,\rm s}$, where $\sigma_{v,0}$ is the gas rms velocity at some large scale, and $\sigma_{v,\rm s}$ is the rms velocity of the stars. However, when a star is formed, its velocity is comparable to that of the nearby gas (except for stars accelerated by dynamical interactions), and then gradually decouples from the gas velocity because of the temporal decorrelation of the turbulence. A turbulent eddy of size $R$, with an rms velocity $\sigma_{v}(R)$, has a turnover time $\tau(R)\sim R/\sigma_{v}(R)$. After that time, the star velocity is decorrelated from the gas velocity at the scale $R$, but remains coupled to the turbulent velocity at larger scales. Thus, the relevant $\sigma_{v}(R)$ for BH infall increases with time, as the star velocity decouples from the gas velocity at increasingly larger scales. The star is also accelerated in the local gravitational potential. Neglecting stars that achieve a significant acceleration by close encounters with other stars (like disrupted binaries), we assume that stars achieve an rms velocity of order the virial velocity at scale $R$, in approximately a dynamical time. According to the two Larson’s relations, $\sigma_{v}(R)$ is also of the order of the virial velocity at the scale $R$, so we adopt a simple approximation where $\sigma_{v,\rm s}=\sigma_{v}$, meaning that both rms velocities grow in time at the same rate, so $\sigma_{v,\rm rel}=\sqrt{2}\,\sigma_{v}(R(t))$, where $R(t)=R(\tau)$, that is we identify the time with the eddy turnover time. Using Larson’s velocity-size and density-size relations from Solomon et al. [42],
@@
Figure 7: Gas-star relative velocity (upper panel) and gas number density (lower panel) versus star age at the same star positions as in Figure 6, but excluding binaries and stars with accretion rate $>5\times 10^{-6}\,M_{\odot}$yr-1 (blue dots). Red squares are rms values (upper panel) and median values (lower panel) computed in age bins, with their least-squares fit, for $t>0.5$ Myr, shown by the red solid lines. The dashed lines are the predictions from Equations 33 and 34, with the velocity normalization increased by a factor 1.6, to match the rms velocity in the simulation.
These relations are hard to test in the simulation, due to the expected scatter, the existence of stars with significant dynamical kicks from close encounters, and the broad range in the extent of the embedded phase preceding the BH phase (or Class II), meaning that the time zero of the decoupling corresponds to a different age for different stars [61, 62]. However, the simulation shows clear evidence of the predicted trends, as illustrated in Figure 7. The figure shows $v_{\rm rel}$ (upper panel) and $n_{\rm H}$ (lower panel) as a function of age (blue dots) for the stars found in the simulation in the same six snapshots used to compute $j_{\rm BH}$. Binaries are not included (35% of the stars) due to their more complex dynamical evolution, nor stars with accretion rate $>5\times 10^{-6}\,M_{\odot}$yr-1 (8% of single stars) because they are still deeply embedded and far from reaching their final mass and starting to decouple from the gas. In the upper panel, we compute the rms relative velocity, $\sigma_{v,\rm rel}$, in age bins (red squares), and fit the result for $t>0.5$ Myr, which gives $\sigma_{v,\rm rel}\sim t^{0.8\pm 0.1}$ (red solid line). The predicted time dependence from Equation 33, with the velocity normalization increased by a factor 1.6 to match the rms velocity in the simulation, is shown by the black dashed line. The lower panel shows the median of $n_{\rm H}$ in age bins (red squares), a least-squares fit for $t>0.5$ Myr giving $n_{\rm H}\sim t^{-2.3\pm 0.4}$ (red solid line), and the prediction from Equation 34 (dashed black line), also based on the renormalized velocity dispersion.
Using Equations 33 and 34, Equation 30 becomes
@@
assuming there is no partial cancellation of angular momentum. Because $\bm{J}$ of the gas captured by an individual star in its trajectory is generally not constant over a $\sim 1$ Myr timescale, there must be some partial cancellation, hence Equations 37 and 38 should be considered as upper limits to the actual PD values.
@@
Acknowledgments
PP and VMP acknowledge financial support by the grant PID2020-115892GB-I00, funded by MCIN/AEI/10.13039/501100011033 and by the grant CEX2019-000918-M funded by MCIN/AEI/10.13039/501100011033. L.P. acknowledges financial support from NSFC under grant No. 11973098 and No. 12373072.

### v2

@@ 1.6 Angular Momentum of Bondi-Hoyle Infall in the Simulation
The simulation can be used also to measure $j_{\rm BH}$ relative to the position and velocity of PMS stars, and to compare the result with the prediction of Equation (4). Although new stars are continuously formed in the simulation, towards the end of the run a significant fraction of them have ages in the approximate range 0.5-2.0 Myr, old enough to be representative of Class II PMS stars. However, because the time it takes to assemble a star may vary from star to star, and can be relatively long ($\sim 1$ Myr) for massive stars [62, 63], we select PMS stars based on the local gas density, rather than the stellar age, which better reflects the observational SED classification as well. For that purpose, we use seven snapshots at regular time intervals covering the last 0.75 Myr of the simulation, yielding a total of 1,629 stellar positions.
@@
Averaging over all seven snapshots, we find that the gas density sampled in spheres of radius $\sim 400$ AU (of the order of the size of the largest observed disks) centered around the stars, $P_{\rm st}(n)$, has a clear bimodal distribution, shown by the shaded blue histogram in the left panel of Supplementary Figure 4. The red unshaded histogram shows the overall gas density distribution sampled uniformly in the whole volume at the same resolution, $P_{\rm V}(n)$. The lower-density peak of $P_{\rm st}(n)$ follows approximately the overall distribution, though shifted to slightly larger density and with a shape a bit skewed to the right, while the higher-density peak has a maximum at a density larger by four orders of magnitude. The stars at such high local density are still embedded in their native dense gas and are generally increasing their mass at a high rate. The black circles in the left panel of Supplementary Figure 4 show the gas infall rate on the sink particles in the simulation (with values shown on the right $y$ axis). It strongly correlates with the local gas density, and the infall rates of stars below $\sim 10^{5}$ cm-3 are of the same order of magnitude as the observed accretion rates of young PMS stars.
@@
We have verified through visualizations that most of these stars at lower densities are accompanied by gas structures with morphology and kinematics consistent with BH trails, and their predicted BH infall rate is also of the order of the measured infall rate in the simulation, as shown in the right panel of Supplementary Figure 4. We have also verified that, on average, the local density is inversely correlated with the stellar age. The stars at low density have clearly decoupled from their native dense cores and tend to sample the random density and velocity fields of the parent cloud at larger scales. Based on these results, we choose a fixed critical density value, $n_{\rm H,cr}$, to select the sink particles representative of Class II PMS stars in the BH phase, $n_{\rm H,cr}=5\times 10^{4}$ cm-3, shown by the vertical dashed line in Figure 6. This selection yields 961 PMS stars, out of the total 1,629 stars found in the seven snapshots.
@@
We measure the mean mass, gas velocity and angular momentum within spheres of different radii centered on the position of each star. The density and velocity fields of the whole 4 pc box are first extracted into a uniform grid of 1,0243 cells, so the cell size is $dx=0.004$ pc or 780 AU. The spheres have radii of 2, 4, 8, 16, 32, 64, 128, and 256 $dx$, with the two smallest and the largest expected to be outside the inertial range of the simulation. The gas mass and the velocity and angular momentum components are averaged within each sphere with a Gaussian cutoff, as in Equation (11).
@@
We estimate the angular momentum of the infalling gas by computing $j$ within a sphere of radius equal to $R_{\rm BH}$, so we need to compute $R_{\rm BH}$ as defined in Equation (3). The mass and velocity of each star are known, as well as the isothermal sound speed in the simulation, $c_{\rm s}=0.18\times 10^{5}$ cm s-1. To compute $v_{\rm rel}$, the difference between the star velocity and the gas velocity, we use the gas mean velocity measured within a sphere centered on the star. Because the flow is turbulent, the mean gas velocity may vary when measured at different scales, so we should compute it at a scale larger than $R_{\rm BH}$, but not too large. We settle on a radius of 8 $dx$=6,250 AU, which is significantly larger than the largest values of $R_{\rm BH}$ and the typical PD sizes. For the sink particles, we find values of $R_{\rm BH}$ in the approximate range 1-3,000 AU. We cannot measure $j$ directly from the simulation at such scales, because they are well within the numerical dissipation range. Instead, we measure it within spheres of radius of 8 $dx$=6,250 AU, as for the estimate of $v_{\rm rel}$, because this is the smallest size we can consider without being significantly affected by numerical dissipation. We then extrapolate the value of $j$ at $R=R_{\rm BH}$, using the linear $j$ scaling established earlier both analytically and numerically (see Supplementary Figure 2).
@@
The results are shown in the right panel of Figure 1 as a function of the stellar mass for all the selected PMS stars in the simulation (blue dots). The least-squares fit has a slope of 0.86 (solid blue line), a bit shallower than the one predicted in Equation (8). The plot shows a significant scatter at any given value of $M_{\rm star}$, as expected from Equation (4) due to the dependence on $\sigma_{v,\rm rel}$: for an individual star at a given time, $j_{\rm BH}$ depends on the local value of $v_{\rm rel}$, which is a random variable. The scatter and the average value of $j$ in PDs are expected to be somewhat smaller than those of individual BH spheres computed here, due to partial cancelation in the vector sums of $\bm{J}$ from different BH spheres. The right panel of Figure 1 shows that observed disks indeed have a slightly reduced mean angular momentum and a smaller scatter than our individual BH spheres, as expected.
@@ 1.7 Time Dependence of PMS Disk Formation by Bondi-Hoyle Infall
Whether or not BH infall is so dominant to completely restructure PDs during the Class II phase, that is on a timescale of $\sim 2$ Myr, depends on the stellar mass and on the time evolution of the relative velocity between the stars and the turbulent gas, $v_{\rm rel}$. The BH infall rate, defined as a mass flux through a surface of area $\pi R_{\rm BH}^{2}$ with gas velocity $v=(c_{\rm s}^{2}+v_{\rm rel}^{2})^{1/2}$, gas number density $n_{\rm H}$, and $R_{\rm BH}$ given by Equation (3) is:
@@
Vorticity corrections to this BH accretion rate [e.g. 64] are not significant in the context of low-mass star formation [65]. When stars are fully decoupled from the parent gas, their velocities are no longer correlated to the gas velocity, and we can assume that the rms velocities satisfy the relation $\sigma^{2}_{v,\rm rel}=\sigma^{2}_{v,0}+\sigma^{2}_{v,\rm s}$, where $\sigma_{v,0}$ is the gas rms velocity at some large scale, and $\sigma_{v,\rm s}$ is the rms velocity of the stars. However, when a star is formed, its velocity is comparable to that of the nearby gas (except for stars accelerated by dynamical interactions) and then gradually decouples from the gas velocity because of the temporal decorrelation of the turbulence. A turbulent eddy of size $R$, with rms velocity $\sigma_{v}(R)$, has a turnover time $\tau(R)\sim R/\sigma_{v}(R)$. After that time, the star velocity is decorrelated from the gas velocity at the scale $R$, but remains coupled to the turbulent velocity at larger scales. Thus, the relevant $\sigma_{v}(R)$ for BH infall increases with time, as the star velocity decouples from the gas velocity at increasingly larger scales. The star is also accelerated in the local gravitational potential. Neglecting stars that achieve significant acceleration by close encounters with other stars (like disrupted binaries), we assume that stars achieve an rms velocity of the order of the virial velocity at scale $R$, in approximately a dynamical time. According to the two Larson’s relations, $\sigma_{v}(R)$ is also of the order of the virial velocity at the scale $R$, so we adopt a simple approximation where $\sigma_{v,\rm s}=\sigma_{v}$, meaning that both rms velocities grow in time at the same rate, so $\sigma_{v,\rm rel}=\sqrt{2}\,\sigma_{v}(R(t))$, where $R(t)=R(\tau)$, that is, we identify the time with the eddy turnover time. Using Larson’s velocity-size and density-size relations from Solomon et al. (1987) [53],
@@
These relations are hard to test in the simulation, due to the expected scatter, the existence of stars with significant dynamical kicks from close encounters, and the broad range in the extent of the embedded phase preceding the BH phase (or Class II), meaning that the time zero of the decoupling corresponds to a different age for different stars [62, 63]. However, the simulation shows clear evidence of the predicted trends, as illustrated in Supplementary Figure 5. The figure shows $v_{\rm rel}$ (upper panel) and $n_{\rm H}$ (lower panel) as a function of age (blue dots) for the stars found in the simulation in the same six snapshots used to compute $j_{\rm BH}$. Binaries are not included (35% of the stars) due to their more complex dynamical evolution, nor stars with accretion rate $>5\times 10^{-6}\,M_{\odot}$yr-1 (8% of single stars) because they are still deeply embedded and far from reaching their final mass and starting to decouple from the gas. In the upper panel, we compute the relative velocity rms, $\sigma_{v,\rm rel}$, in age bins (red squares), and fit the result for $t>0.5$ Myr, which gives $\sigma_{v,\rm rel}\sim t^{0.8\pm 0.1}$ (solid red line). The predicted time dependence from Equation (33), with the velocity normalization increased by a factor 1.6 to match the rms velocity in the simulation, is shown by the black dashed line. The lower panel shows the median of $n_{\rm H}$ in age bins (red squares), a least-squares fit for $t>0.5$ Myr giving $n_{\rm H}\sim t^{-2.3\pm 0.4}$ (red solid line), and the prediction from Equation (34) (dashed black line), also based on the renormalized velocity dispersion.
Using Equations (33) and (34), Equation (30) becomes
@@
assuming there is no partial cancellation of angular momentum. Because $\bm{J}$ of the gas captured by an individual star in its trajectory is generally not constant over a $\sim 1$ Myr timescale, there must be some partial cancellation, hence Equations (37) and (38) should be considered as upper limits to the actual PD values.
2 Supplementary Figures
Figure 1: Correlation and structure function scalings. Left: Longitudinal (red squares) and transverse (blue circles) components of the mixed correlation function $B^{\rho,v}_{ij}=\langle\rho_{1}\rho_{2}v_{1i}v_{2j}\rangle$. Under the assumption of independence between $\rho$ and ${\bm{v}}$, both are expected to equal $v^{\prime 2}B_{\rho}$ (black circles). The density correlation function, $B_{\rho}$, exhibits a power-law scaling with an exponent of 0.61 in the inertial range. Right: Longitudinal (circles) and transverse (squares) components the velocity structure tensor with (red) and without (blue) density weighting. The structure functions show similar scalings, but the amplitudes of the density-weighted ones are slightly smaller.
Figure 2: Scaling of the variance of the specific angular momentum, $\langle j^{2}\rangle^{1/2}$, as a function of the size $R$ in units of half the box size, $L$. The blue square symbols and the connected blue lines are the results averaged over four snapshots of the simulation. The red dashed line is the linear prediction from Equation (21).
Figure 3: Scaling relations of the specific angular momentum. The specific gas angular momentum within the BH radius, $j_{\rm BH}$, versus stellar mass, $M_{\rm star}$, measured in 7 snapshots of the simulation, for all sink particles identified as accreting PMS stars, is shown by the blue dots, with the least-squares fit shown by the blue solid line. The two dashed lines correspond to $j_{\rm BH}$ predicted in Equation (8), for $t=1$ Myr (upper line) and 4 Myr (lower line). The dashed-dotted lines correspond to the disk specific angular momentum, $j_{\rm d}$, for given disk radii, as in Equation (29). The red empty circles give the observational values of $M_{\rm star}$ and $j_{\rm d}$ for PMS stars with resolved disk sizes [33, 35, 42, 43], with the least-squares fit shown by the red solid line. This figure is the same as the right panel of Figure 1, but with the $j$ values of the observed PMS disks derived from the size of the gas disk, $R_{\rm d}=R_{\rm CO}$, instead of the dust disk.
Figure 4: Infall rates and local gas density around the stars in the simulation. Left: Probability distribution of density sampled in spheres of radius $\sim 403$ au centered around the stars, $P_{\rm st}(n_{\rm H})$ (blue shaded histogram), and sampled uniformly in the whole volume at the same resolution, $P_{\rm V}(n_{\rm H})$ (red unshaded histogram), using all seven time snapshots of the simulation. The black points give the infall rate on the stars from the simulation, averaged over a period of 5,000 yr, with values shown in the right $y$ axis and a linear least-squares fit shown by the long dashed line. The vertical dashed line corresponds to the critical density, $n_{\rm H,cr}=5\times 10^{4}$ cm-3, used to select the stars representative of Class II objects in the BH phase. Right: The predicted BH infall rate versus the infall rate of the stars in the simulation (as in the left plot). The red symbols are the Class II stars based on the critical density criterion, and the red line is the least-squares fit to the red symbols. The black line is the one-to-one relation.
Figure 5: Time evolution of the star-gas decoupling. Gas-star relative velocity (upper panel) and gas number density (lower panel) versus star age at the same star positions as in Supplementary Figure 4, but excluding binaries and stars with accretion rate $>5\times 10^{-6}\,M_{\odot}$yr-1 (blue dots). Red squares are rms values (upper panel) and median values (lower panel) computed in age bins, with their least-squares fit, for $t>0.5$ Myr, shown by the red solid lines. The dashed lines are the predictions from Equations (21) and (22), with the velocity normalization increased by a factor 1.6, to match the rms velocity in the simulation.
@@ 2 Supplementary Figures
Carbon Footprint The total cost of the simulation used in this work (and in 7 prior publications where the CO2e was not stated) was approximately 50M core hours (Xeon E5-2680v3), corresponding to 847MWh and a carbon footprint of 121t CO2e (calculated using green-algorithms.org v2.2 [66]).
[…]
Data Availability Supplemental material can be obtained from a dedicated public URL (http://www.erda.dk/vgrid/BH-Accretion-of-Disks). This includes tabulated data to reproduce the plots in the paper, and movies of the time evolution of BH trails for some of the stars in the simulation. Raw and post-processed data from the simulation are available upon request (email requests to Paolo Padoan, ppadoan@icc.ub.edu).
[…]
Code Availability The MHD and gravity solvers used in this study are closely related to the public RAMSES version available at https://bitbucket.org/rteyssie/ramses. The numerical methods, developed in Copenhagen, relevant for star formation are described in Haugbølle et al. (2018) [57].
[…]
Acknowledgments We thank the anonymous referees for helpful comments. P.P. acknowledges support from the grants PID2020-115892GB-I00 and CEX2019-000918-M funded by MCIN/AEI/10.13039/501100011033, and by the US National Science Foundation under Grant AST 2408023. L.P. acknowledges support by the NSFC under grants No. 11973098 and No. 12373072. T.H. acknowledges support from the Independent Research Fund Denmark through grant No. DFF 802100350B. VM.P. acknowledges financial support by the grants PID2020-115892GB-I00 and CEX2019-000918-M funded by MCIN/AEI/10.13039/501100011033, and by the European Research Council via the ERC Synergy Grant “ECOGAL” (project ID 855130). The Tycho supercomputer hosted at the SCIENCE HPC center at the University of Copenhagen was used for carrying out the postprocessing, analysis, and long-term storage of the results.
[…]
Author Contributions Statement P.P. and VM.P. conceived, designed, and carried out the data-analysis study, and L.P. the analytical modeling. T.H. optimized the code and ran the simulation. Å.N. contributed to the code optimization and conceived with P.P. the astrophysical scenario. P.P. and L.P. co-wrote the paper and all other authors contributed edits to the manuscript.
[…]
Competing Interests Statement The authors declare no competing interests.